OpenStudy (anonymous):

1/2 ft = how many inches?

OpenStudy (anonymous):

6 inches

OpenStudy (anonymous):

1ft 12in --- = ---- 1/2 in? cross mutipfly [(1/2)*12]/1=6in If I deserve medal then give me one

OpenStudy (anonymous):

amu, seriously? lol

OpenStudy (anonymous):

a i said i ws just checking if these r right!!!

OpenStudy (anonymous):

wait....1 yard minus 9 inches is???

OpenStudy (anonymous):

1yard = 36in 36in -9in = 27 in

OpenStudy (anonymous):

24 oz + how many cups?

OpenStudy (anonymous):

3 cups

OpenStudy (anonymous):

1 feet =12 inch so, half feet=6 inch.

OpenStudy (anonymous):

medel plz... m giving u alll :)

OpenStudy (anonymous):

oops i meant = sign not the+ sign

OpenStudy (anonymous):

??

OpenStudy (anonymous):

0.5divide by 1/2

OpenStudy (anonymous):

.5=1/2 so (1/2)/(1/2)=1

OpenStudy (anonymous):

n what do u want?